# Ask, and it shall be given: Turing completeness of prompting

### v1

@@ 3.4 Construction of Transformer
In this subsection, we sketch how to construct a decoder-only Transformer $\varGamma$ that executes prompts through CoT steps as described in Section 3.2. This means that we need to execute $2$-PTMs using ReLU activation, layer normalization, and causal attention. Due to the space limit, we only present three core operations to be used by $\varGamma$ here and defer the detailed construction to Appendix B.
Boolean algebra via ReLU activation.
Let $\operatorname{ReLU}(z):=\max\{z,0\}$ denote the ReLU function. Boolean algebra is a basic building block of our Transformer for, e.g., checking the go-to condition. A core operation in Boolean algebra is the $\land$ operation. Here, we implement $\land$ via ReLU activation:
@@
Equality check via layer normalization.
Let $\operatorname{LN}(\boldsymbol{z}):=\frac{\boldsymbol{z}}{\|\boldsymbol{z}\|_{2}}1_{[\boldsymbol{z}\neq\boldsymbol{0}]}$ denote layer normalization (LN). When checking whether a tape cell has been written or not, we will need an equality check between two real numbers: $\operatorname{Equal}:\mathbb{R}^{2}\to\{0,1\}$ where $\operatorname{Equal}(u,v):=1_{[u=v]}$ ($u,v\in\mathbb{R}$). Since $\operatorname{Equal}$ is not a continuous map, it cannot be implemented using only affine maps or ReLU activation. Instead, we implement $\operatorname{Equal}$ via layer normalization as follows:
@@
Farthest retrieval via causal attention.
A core operation to be used by the Transformer $\varGamma$ can be abstracted as follows: Given a sequence of values $\boldsymbol{v}=\langle v_{0},\ldots,v_{|\boldsymbol{v}|-1}\rangle$ with $v_{i}\in\{-1,0,+1\}$ for all $i$, find the smallest $i$ such that $\sum_{j=0}^{i}v_{j}=\sum_{j=0}^{|\boldsymbol{v}|-1}v_{j}$. However, since causal attention is an average rather than a sum, it cannot even compute the sum $\sum_{j=0}^{i}v_{j}$ for any $i$. It can compute the average $\frac{1}{i+1}\sum_{j=0}^{i}v_{j}$ for every $i$, but $\sum_{j=0}^{i}v_{j}=\sum_{j=0}^{|\boldsymbol{v}|-1}v_{j}$ does not mean $\frac{1}{i+1}\sum_{j=0}^{i}v_{j}=\frac{1}{|\boldsymbol{v}|}\sum_{j=0}^{|\boldsymbol{v}|-1}v_{j}$. To bypass the mismatched coefficients $\frac{1}{i+1}$ and $\frac{1}{|\boldsymbol{v}|}$, we use LN to remove the averaging coefficient $\frac{1}{i+1}$. First, we let every token attend to the initial token ^ at $i=0$ to compute $\frac{1}{i+1}$ and compute
@@ 4 Complexity bounds
In this section, we (i) show in Section 4.1 that our $2$-PTMs are Turing-complete and nearly as efficient as TMs and (ii) use this result to characterize the complexities of our constructed $\varGamma$ in Sections 4.2 & 4.3. Throughout this section, we let $t(n)$ denote a complexity function. Following the convention in complexity theory, we allow different computable functions to have different constant factors in the big $\mathrm{O}$ notation.
@@ 5 Conclusion
Our current work focuses on expressive power rather than learnability. While we have shown the existence of a Transformer on which prompting is Turing-complete, it does not necessarily imply that a Transformer effectively learns to simulate any $2$-PTMs through CoT steps. Investigating the learnability of such a Transformer is an intriguing direction for future research.
@@ A.2 Bi-infinite two-tape TMs simulating $2$-PTMs
If $\iota_{j}=\tau\texttt{!}_{k}$ ($\tau\in\{\mathtt{A},\mathtt{B}\}$), we go to $k$ if the pointed cell on tape $\tau$ is $\mathtt{1}$ and to $j+1$ otherwise:

### v2

@@ 3.4 Construction of Transformer
In this subsection, we sketch how to construct a decoder-only Transformer $\varGamma$ that executes prompts through CoT steps as described in Section 3.2. Due to the space limit, we only present three core operations to be used by $\varGamma$ here and defer the detailed construction to Appendix B.
Boolean algebra via ReLU activation. Let $\operatorname{ReLU}(z):=\max\{z,0\}$ denote the ReLU function. Boolean algebra is a basic building block of our Transformer for, e.g., checking the go-to condition. A core operation in Boolean algebra is the $\land$ operation. Here, we implement $\land$ via ReLU activation:
@@
Equality check via layer normalization. Let $\operatorname{LN}(\boldsymbol{z}):=\frac{\boldsymbol{z}}{\|\boldsymbol{z}\|_{2}}1_{[\boldsymbol{z}\neq\boldsymbol{0}]}$ denote layer normalization (LN). When checking whether a tape cell has been written or not, we will need an equality check between two real numbers: $\operatorname{Equal}:\mathbb{R}^{2}\to\{0,1\}$ where $\operatorname{Equal}(u,v):=1_{[u=v]}$ ($u,v\in\mathbb{R}$). Since $\operatorname{Equal}$ is not a continuous map, it cannot be implemented using only affine maps or ReLU activation. Instead, we implement $\operatorname{Equal}$ via layer normalization as follows:
@@
Farthest retrieval via causal attention. A core operation to be used by the Transformer $\varGamma$ can be abstracted as follows: Given a sequence of values $\boldsymbol{v}=\langle v_{0},\ldots,v_{|\boldsymbol{v}|-1}\rangle$ with $v_{i}\in\{-1,0,+1\}$ for all $i$, find the smallest $i$ such that $\sum_{j=0}^{i}v_{j}=\sum_{j=0}^{|\boldsymbol{v}|-1}v_{j}$. However, since causal attention is an average rather than a sum, it cannot even compute the sum $\sum_{j=0}^{i}v_{j}$ for any $i$. It can compute the average $\frac{1}{i+1}\sum_{j=0}^{i}v_{j}$ for every $i$, but $\sum_{j=0}^{i}v_{j}=\sum_{j=0}^{|\boldsymbol{v}|-1}v_{j}$ does not mean $\frac{1}{i+1}\sum_{j=0}^{i}v_{j}=\frac{1}{|\boldsymbol{v}|}\sum_{j=0}^{|\boldsymbol{v}|-1}v_{j}$. To bypass the mismatched coefficients $\frac{1}{i+1}$ and $\frac{1}{|\boldsymbol{v}|}$, we use LN to remove the averaging coefficient $\frac{1}{i+1}$. First, we let every token attend to the initial token ^ at $i=0$ to compute $\frac{1}{i+1}$ and compute
@@ 4 Complexity bounds
We (i) show in Section 4.1 that $2$-PTMs are Turing-complete and nearly as efficient as TMs and (ii) use this result to characterize the complexities of our constructed $\varGamma$ in Sections 4.2 & 4.3. Throughout this section, let $t(n)$ denote a complexity function. Following the convention in complexity theory, we allow different computable functions to have different constant factors in big $\mathrm{O}$.
@@ A.2 Bi-infinite two-tape TMs simulating $2$-PTMs
If $\iota_{j}=\tau\texttt{!}_{k}$ ($\tau\in\{\mathtt{A},\mathtt{B}\}$), we go to $k$ if the pointed cell on tape $\tau$ is $\mathtt{0}$ and to $j+1$ otherwise:
@@ B.9 Generating next token
B.10 Composing the Transformer
In this subsection, we describe how to compose the aforementioned operations into a Transformer.
Embedding layer. Recall that the construction uses 4 positional encodings $p_{i},z_{i}^{\text{pos, 1}},z_{i}^{\text{pos, 2}},z_{i}^{\text{pos, 3}}$. Thus, our token embedding and positional encoding use the first $|\varSigma|+1$ dimensions. For the token embedding $\operatorname{emb}$, the first $|\varSigma|$ dimensions are the ont-hot representation of the token, and the next dimension is zero. For example, if $v_{i}$ is the first token in $\varSigma$, then
$\operatorname{emb}(v_{i})=(\underbrace{1,0,\dots,0,}_{\text{first}|\varSigma|\text{ dims: one-hot}}\quad\underbrace{0,}_{\text{next dim: zero}}\quad\underbrace{0,0,\ldots,0}_{\text{slots for intermediate results}})^{\mathsf{T}}.$ (121)
For the positional encoding $\operatorname{pos}$,
$\operatorname{pos}(i)=(\underbrace{0,0,\dots,0,}_{\text{first }|\varSigma|\text{ dims: zeros}}\quad\underbrace{p_{i},}_{\text{next dim: pos enc}}\underbrace{0,0,\ldots,0}_{\text{slots for intermediate results}})^{\mathsf{T}}.$ (122)
Therefore, the embedding layer for the example above is
$\operatorname{emb}(v_{i})+\operatorname{pos}(i)=(\underbrace{1,0,\dots,0,}_{\text{first }|\varSigma|\text{ dims: one-hot}}\quad\underbrace{p_{i},}_{\text{next dim: pos enc}}\quad\underbrace{0,0,\ldots,0}_{\text{slots for intermediate results}})^{\mathsf{T}}.$ (123)
Here, $\operatorname{emb}(v_{i})+\operatorname{pos}(i)$ is indeed the standard embedding layer in LLMs.
Intermediate results. Recall that each Transformer layer has a residual connection after its output. Thus, we can compute each intermediate result via a Transformer layer and add it to the hidden embedding via the residual connection. For the example above, we can construct a Transformer layer that computes $z_{i}^{\text{pos, 1}}$, and then the hidden embedding after the residual connection of this Transformer layer is
$(\underbrace{1,0,\dots,0,}_{\text{first }|\varSigma|\text{ dims: one-hot}}p_{i},z_{i}^{\text{pos, 1}},\underbrace{0,\ldots,0}_{\text{slots for other intermediate results}})^{\mathsf{T}}.$ (124)
B.11 Size of the constructed Transformer
In this subsection, we show that the constructed Transformer $\varGamma$ has a constant size in terms of the number of its operations and the number of bits to represent its parameters.
Number of operations. From the construction above, we can see that the constructed Transformer $\varGamma$ does not depend on $\varphi$ or $\boldsymbol{x}$. (To compute different functions $\varphi$, we change only the prompt but do not need to change the Transformer.) It is clear to see that our constructed Transformer has a constant number of operations.
Number of parameter bits. From the construction above, we can see that the constructed Transformer uses only $30,\frac{1}{2},1,2,$ 3 as its parameters. These numbers can be exactly expressed with a constant number of bits.
Appendix C Experiments
We demonstrate our construction through a Python implementation. Our code and demo results are publicly available at https://github.com/q-rz/ICLR25-prompting-theory.
Appendix D Concluding remarks
Connection with universal TMs.
Hennie & Stearns (1966) have shown that there exists a universal TM (UTM) that can simulate any Turing machine $M$ in $\mathrm{O}(t(n)\log t(n))$ steps if $M$ halts in $t(n)\geq n$ steps. Then by Pérez et al. (2019), there exists a Transformer that simulates this UTM. However, due to the complication of the UTM, it would be quite involved to explicitly construct this Transformer. Nevertheless, this UTM has inspired us to propose $2$-PTMs, which enables us to explicitly construct a simpler Transformer. Furthermore, our $2$-PTMs establish the first theoretical framework to formalize prompting. Using our $2$-PTM-based framework, we believe that people will be able to generalize more results from the classic one-model-one-task paradigm to the LLM prompting paradigm.
Discussion on CoT complexity. A limitation of this work is that our construction uses long CoTs for hard problems with high time complexity. However, while it might be possible to slightly improve the CoT complexity, recent theoretical evidence has suggested that long CoTs are very likely to be necessary for these hard problems. For example, Merrill & Sabharwal (2024a) have shown that any Transformer with $\mathrm{O}(\log n)$ CoT steps can solve only $\mathsf{L}$ problems. Assuming $\mathsf{L}\neq\mathsf{NL}$, it implies that any Transformer with $\mathrm{O}(\log n)$ CoT steps cannot even solve any $\mathsf{NL}$-complete problems such as directed graph connectivity. An interesting future work is to show a tight lower bound of the CoT complexity.
Discussion on hardmax. Following prior works (e.g., Pérez et al., 2019; Hahn, 2020), this work uses $\operatorname{hardmax}$ in attention to simplify construction. However, real-world Transformers use $\operatorname{softmax}$ in attention. Using $\operatorname{hardmax}$ instead of $\operatorname{softmax}$ is a limitation of this area. We hope to address this limitation in future work.
Discussion on context windows. The Transformer constructed in this work uses an unbounded context window. However, practical implementations of Transformers typically use a bounded context window due to memory considerations. This is indeed a limitation of current studies on the Turing completeness of Transformers. To address this limitation, we hope to study bounded context windows in future work.
Discussion on learnability. Our current work focuses on expressive power rather than learnability. While we have shown the existence of a Transformer on which prompting is Turing-complete, it does not necessarily imply that a Transformer effectively learns to simulate any $2$-PTMs through CoT steps. Investigating the learnability of such a Transformer is an intriguing direction for future research.